OpenStudy (anonymous):

What is the equation for the pair of points (-5, -3) and (-1,-9)

OpenStudy (kaiz122):

y - y1 = m(x - x1), where m(slope) =\[\frac{y_{2}-y_{1}}{x_2-x_1} \]

OpenStudy (anonymous):

I have that formula, but I was unable to come up with the correct answer. I'm very confused, and this is my last problem!

OpenStudy (kaiz122):